OpenStudy (anonymous):

the perimeter of a rectangle is 106m the length is 5m more thatn twice the width.what would be the demensions

OpenStudy (anonymous):

2l+2w=106 2w=l+5

OpenStudy (anonymous):

2l+2w=106 -l+2w=5

OpenStudy (anonymous):

"length is 5 more than twice the width" shouldn't that be L=2W+5

OpenStudy (anonymous):

yes my bad

OpenStudy (anonymous):

it is very late here also

OpenStudy (anonymous):

anyways 2l+2w=106 l-2w=5

OpenStudy (anonymous):

3l = 111

OpenStudy (anonymous):

2(x+5) + 2x = 106 4x = 96 x = 24 ----> width x+5 = 299 --> length

OpenStudy (anonymous):

299 can not be a length

OpenStudy (anonymous):

oopps, its 29,

OpenStudy (anonymous):

your second statement becomes false with those answers

OpenStudy (anonymous):

2w+5 does not equal 29

OpenStudy (anonymous):

l=37 w=16

OpenStudy (anonymous):

dhashni is right i believe... i got 37 for one before i had to chime in

OpenStudy (anonymous):

you forgot a 2 in your case it should have been 2(2x+5)+2x= 106

OpenStudy (anonymous):

@nikki, here is the correct solution 2L+2W=106 L=2W+5 Subst L from equ 2 into equ 1 2(2W+5)+2W=106 4W+10+2W=106 6W=96 W=16 Now sub 16 into W in equ 2 L=2(16)+5=32+5=37 Length=37; width=16